OpenStudy (anonymous):

Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $123.00 for 3 days and 300 miles, while Mary was charged $216.00 for 5 days and 600 miles. What does Best Rental charge per day, for mileage?

OpenStudy (anonymous):

Let the daily fee be d and the mileage fee be m.\[3d+300m=123\]Thus\[6d+600m=246\] And \[5d+600m=216\] Thus (subtracting) \[d=30\]And from the first equation \[m=0.11\]